# LiNo: Advancing Recursive Residual Decomposition of Linear and Nonlinear Patterns for Robust Time Series Forecasting

### v1

@@ D.1 Visualization of LiNo’s Forecasting Results
To provide insights that help readers better understand the working mechanism of LiNo and to intuitively grasp the effects of advanced ‘Recursive Residual Decomposition’ (RRD), we present visualizations of LiNo’s forecasting results (Last channel/variate) across 13 multivariate forecasting benchmarks. We set the input length $T=96$, prediction length $F=96$, and number of LiNo blocks (layers) $N=3$. ‘LP’ denotes Linear prediction, and ‘NP’ stands for Nonlinear prediction. LP $i$ or NP $i$ ($i\in\{1,2,3\}$) is the linear or nonlinear prediction of $i$-th layer (level). Results are in Figure 6 18.
D.2 Visualization of Weight of Li blocks and No blocks
@@
We present visualizations of the weight of Li blocks and No blocks obtained across 13 multivariate forecasting benchmarks to help better understand the proposed LiNo. We set the input length $T=96$, prediction length $F=96$, and number of LiNo blocks (layers) $N=3$. Results are in Figure 19 31.
@@
Appendix E Full Results
E.1 Full Results of Multivariate Forecasting Benchmark
@@
E.2 Full Results of Univariate Forecasting Benchmark

### v4

@@ G.1 Visualization of LiNo’s Forecasting Results
To provide insights that help readers better understand the working mechanism of LiNo and to intuitively grasp the effects of advanced ‘Recursive Residual Decomposition’ (RRD), we present visualizations of LiNo’s forecasting results (Last channel/variate) across 13 multivariate forecasting benchmarks. We set the input length $T=96$, prediction length $F=96$, and number of LiNo blocks (layers) $N=3$. ‘LP’ denotes Linear prediction, and ‘NP’ stands for Nonlinear prediction. LP $i$ or NP $i$ ($i\in\{1,2,3\}$) is the linear or nonlinear prediction of $i$-th layer (level). Results are in Figure 5 17.
G.2 Visualization of Weight of Li blocks and No blocks
@@
We present visualizations of the weight of Li blocks and No blocks obtained across 13 multivariate forecasting benchmarks to help better understand the proposed LiNo. We set the input length $T=96$, prediction length $F=96$, and number of LiNo blocks (layers) $N=3$. Results are in Figure 18 30.
@@
Appendix H Full Results
H.1 Full Results of Multivariate Forecasting Benchmark
@@
H.2 Full Results of Univariate Forecasting Benchmark